# ValveFit: An analysis-suitable B-spline-based surface fitting framework for patient-specific modeling of tricuspid valves

### v1

@@ Abstract
Patient-specific computational modeling of the tricuspid valve (TV) is vital for the clinical assessment of heart valve diseases. However, this process is hindered by limitations inherent in the medical image data, such as noise and sparsity, as well as by complex valve dynamics. We present ValveFit, a novel GPU-accelerated and differentiable B-spline surface fitting framework that enables rapid reconstruction of smooth, analysis-suitable geometry from point clouds obtained via medical image segmentation. We start with an idealized TV B-spline template surface and optimize its control point positions to fit segmented point clouds via an innovative loss function, balancing shape fidelity and mesh regularization. Novel regularization terms are introduced to ensure that the surface remains smooth, regular, and intersection-free during large deformations. We demonstrate the robustness and validate the accuracy of the framework by first applying it to simulation-derived point clouds that serve as the ground truth. We further show its robustness across different point cloud densities and noise levels. Finally, we demonstrate the performance of the framework toward fitting point clouds obtained from real patients at different stages of valve motion. An isogeometric biomechanical valve simulation is then performed on the fitted surfaces to show their direct applicability toward analysis. ValveFit enables automated patient-specific modeling with minimal manual intervention, paving the way for the future development of direct image-to-analysis platforms for clinical applications.
@@ 1 Introduction
In the past decade, advancements in non-invasive cardiac imaging methods have significantly improved the clinical diagnosis and treatment of heart valve diseases [13]. These advanced imaging modalities enable clinicians to generate time-resolved three-dimensional images non-invasively, allowing for preemptive and personalized treatment planning. In cardiac imaging, spatiotemporal echocardiography (4D echo) has emerged as an established non-invasive modality for the visualization of the heart valves, serving as an evaluation of the valve structure and function by providing anatomical and functional information [14]. Despite these advances in non-invasive clinical imaging that enable accurate and earlier prediction of the onset of valve regurgitation and personalized treatment planning [15, 16], technical challenges persist in patient-specific computational modeling of the heart valves [17]. The inherent noise, sparsity, and low resolution of imaging modalities such as echocardiography limit the accurate segmentation and visualization of tricuspid heart valves. Addressing these challenges requires conversion of this low-fidelity imaging data to high-fidelity geometric representations that offer high-resolution surface visualization [18]. The solution also necessitates the integration of advanced surface fitting methods that can facilitate the automatic generation of patient-specific models, which can be directly used for computational biomechanical analysis [19].
In this paper, we present a novel surface fitting framework that automatically generates patient-specific spline-based surface geometry of the tricuspid valve. The proposed method performs smooth and regular surface deformation of a template surface to fit the 3D point cloud obtained from echocardiographic image segmentation. We validate our framework using tricuspid valve surfaces obtained from simulation data, showing robustness toward different point cloud sampling and noise intensities. We then demonstrate the fitting results using point cloud data from real patients and the application of isogeometric analysis on the fitted surfaces  [20]. The rest of the article is organized as follows. In Section 2, we provide an overview of patient-specific computational modeling frameworks for heart valves. The proposed ValveFit framework is introduced in Section 3 and the surface fitting results along with a discussion on the performance are provided in Section 4. We also demonstrate that the surfaces generated by ValveFit are analysis-suitable by performing a simplified valve closure simulation. Finally, we provide concluding remarks and outline future directions in Section 5.
@@ 2 Related Work on Patient-Specific Modeling of Heart Valves
Heart valve surface fitting methods play a crucial role in the functional and structural assessment of the valves by providing accurate geometric representations from medical imaging data [25, 26, 27]. Several frameworks have been presented in the realm of patient-specific modeling based on surface fitting of point clouds obtained from image segmentation [28, 29]. For example, least squares optimization has been used to fit surfaces to point cloud segmentation data for the tricuspid valves [30]. These methods rely on good-quality image segmentation data to generate analysis-suitable meshes. If the point cloud data from images is sparse, noisy, and incomplete, these existing methods will likely overfit the data, resulting in inaccurate patient-specific valve models that would require considerable manual intervention to fix the models in order to achieve good mesh quality for analysis [31, 32]. Biomechanical models have been used to accurately fit the physiology of the heart valves [8, 33, 34, 35]. However, these methods do not impose stricter geometric regularization constraints to ensure good-quality surfaces during fast valve motion. On the other hand, machine learning-based approaches have been used to model the heart valve geometry directly from medical images, circumventing the image segmentation procedure [36]. However, the accuracy of these methods is dependent on large data from diverse populations, which is typically not available.
@@
  • A GPU accelerated B-spline surface fitting framework using automatic differentiation, enabling faster B-spline surface evaluations and gradient computation.
@@
3 Methods
In this section, we provide a detailed description of the proposed valve fitting framework, ValveFit. We first describe the image acquisition process and the segmentation of the tricuspid valve geometry from imaging data in Section 3.1. We then describe the proposed surface fitting framework for the automatic generation of patient-specific valve geometries in Section 3.2.
3.1 Image Acquisition and Segmentation
@@ 3.2.1 The Template Geometry
3.2.2 The Loss Function
The proposed ValveFit framework performs iterative deformation through the minimization of a loss function using gradient descent optimization. To ensure the generation of an analysis-suitable, patient-specific TV geometry [72, 73], we enforce smooth and regular deformation of the template surface at each iteration. We define the loss function ($\mathcal{L}$), consisting of a shape fidelity term and a mesh regularization term, as
@@ 3.2.2 The Loss Function
where $\mathbf{S}$ denotes the template B-spline surface and $\mathbf{Q}$ denotes the target point cloud. The shape fidelity term $\mathcal{E}_{\text{fid}}(\mathbf{S},\mathbf{Q})$ ensures that the fitted surface closely approximates the target point cloud. It is defined as
$\mathcal{E}_{\text{fid}}(\mathbf{S},\mathbf{Q})=w_{\text{cd}}\,d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})+w_{\text{hd}}\,d_{\mathrm{HD}}(\mathbf{S},\mathbf{Q})+w_{\text{a}}\,d_{\mathrm{CD}}(\mathbf{A}_{\text{S}},\mathbf{A}_{\text{Q}})\text{ ,}$ (3)
where $d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})$ denotes the one-sided Chamfer distance. $d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})$ evaluates the average squared distance between every point in the point cloud and the closest point on the template surface. The Hausdorff distance, $d_{\mathrm{HD}}(\mathbf{S},\mathbf{Q})$, measures the maximum distance between the template surface and the point cloud. $d_{\mathrm{CD}}(\mathbf{A}_{\text{S}},\mathbf{A}_{\text{Q}})$ calculates the Chamfer distance between the parametric boundary on the template surface closest to the annulus $\mathbf{A}_{\text{S}}:=\mathbf{S}(u=u_{0})$ and the points corresponding to the annulus in the target point cloud $\mathbf{A}_{\text{Q}}$. $w_{\text{cd}}$, $w_{\text{hd}}$, and $w_{\text{a}}$ are the weights controlling the relative contributions of each distance term within the shape fidelity term.
In Eq. (3), $d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})$, $d_{\mathrm{HD}}(\mathbf{S},\mathbf{Q})$ and $d_{\mathrm{CD}}(\mathbf{A}_{\text{S}},\mathbf{A}_{\text{Q}})$ are computed using a distance matrix, $\mathbf{D}\in\mathbb{R}^{N\times M}$, which encodes pairwise distances between the points sampled on the template surface and those in the target point cloud. Let $\mathbf{s}_{k}=\mathbf{S}(u_{k},v_{k})$, for $k=1,...,M$, denote points sampled from the template surface $\mathbf{S}$, and let $\mathbf{q}_{l}\in\mathbf{Q}$, for $l=1,...,N$, denote points in the target point cloud $\mathbf{Q}$. Each element $D_{kl}$ of the distance matrix is defined as
$D_{kl}=||\mathbf{s}_{k}-\mathbf{q}_{l}||,\quad\forall\,\mathbf{s}_{k}\in\mathbf{S}\,,\,\,\mathbf{q}_{l}\in\mathbf{Q}\text{ .}$ (4)
@@
$d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})=\frac{1}{N}\sum_{l=1}^{N}m^{Q}_{l}.$ (5)
Next, the Hausdorff distance is defined as
$d_{\mathrm{HD}}(\mathbf{S},\mathbf{Q})=\max\{\max_{k}\,m^{S}_{k}\,;\,\max_{l}\,m^{Q}_{l}\}.$ (6)
The Hausdorff distance term ensures that the largest deviations between the surface and the point cloud are minimized, hence addressing outliers in the point cloud, particularly at the boundaries. Finally, the computation of $d_{\text{CD}}(\mathbf{A}_{S},\mathbf{A}_{Q})$ follows an approach similar to that used for computing $d_{\text{CD}}$. However, it computes the two-sided Chamfer distance, which also includes the mean of the nearest neighbor distances from $\mathbf{A}_{S}$ to $\mathbf{A}_{Q}$. The sensitivity to outliers and noise can be modulated by adjusting the weights $w_{\text{cd}}$, $w_{\text{hd}}$, and $w_{\text{a}}$ in Eq. (3).
@@
The regularization term $\mathcal{E}_{\text{reg}}(\mathbf{S})$ imposes smoothness and regularity constraints on the surface to obtain a smooth parameterization in the physical domain and prevent element distortions or self-intersections. It comprises two terms:
$\mathcal{E}_{\text{reg}}(\mathbf{S})=w_{\text{orth}}\,R_{\text{orth}}(\mathbf{S})+w_{\text{norm}}\,R_{\text{norm}}(\mathbf{S}).$ (7)
@@
The weights $w_{\text{orth}}$ and $w_{\text{norm}}$ allow us to control the influence of each of the above two terms in Eq. (7), allowing for adjustments based on specific mesh quality requirements. We can efficiently compute the tangent vectors and normal vectors required for Eq. 7 by using the automatic differentiation capabilities of JAX [74]. We evaluate the partial derivatives of the surface coordinates ($x$, $y$, and $z$) with respect to the parametric coordinates $u$ and $v$ by evaluating the Jacobian matrix, $J$, using automatic differentiation as

### v2

@@ Abstract
Patient-specific computational modeling of the tricuspid valve (TV) is vital for the clinical assessment of heart valve diseases. However, this process is hindered by limitations inherent in the medical image data, such as noise and sparsity, as well as by complex valve dynamics. We present ValveFit, a novel differentiable B-spline surface fitting framework that enables rapid reconstruction of smooth, analysis-suitable geometry from point clouds obtained via medical image segmentation. We start with an idealized TV B-spline template surface and optimize its control point positions to fit segmented point clouds via an innovative loss function, balancing shape fidelity and mesh regularization. Novel regularization terms are introduced to ensure that the surface remains smooth, regular, and intersection-free during large deformations. We demonstrate the robustness and validate the accuracy of the framework by first applying it to simulation-derived point clouds that serve as the ground truth. We further show its robustness across different point cloud densities and noise levels. Finally, we demonstrate the performance of the framework toward fitting point clouds obtained from real patients at different stages of valve motion. An isogeometric biomechanical valve simulation is then performed on the fitted surfaces to show their direct applicability toward analysis. ValveFit enables automated patient-specific modeling with minimal manual intervention, paving the way for the future development of direct image-to-analysis platforms for clinical applications.
@@ 1 Introduction
In the past decade, advancements in non-invasive cardiac imaging methods have significantly improved the clinical diagnosis and treatment of heart valve diseases [13]. These advanced imaging modalities enable clinicians to generate time-resolved three-dimensional images non-invasively, allowing for preemptive and personalized treatment planning. In cardiac imaging, spatiotemporal echocardiography (4D echo) has emerged as an established non-invasive modality for the visualization of the heart valves, serving as an evaluation of the valve structure and function by providing anatomical and functional information [14]. Despite these advances in non-invasive clinical imaging that enable accurate and earlier prediction of the onset of valve regurgitation and personalized treatment planning [15, 16], technical challenges persist in patient-specific computational modeling of the heart valves [17]. The inherent noise, sparsity, and low resolution of imaging modalities such as echocardiography limit the accurate segmentation and visualization of tricuspid heart valves. Addressing these challenges requires the conversion of this low-fidelity imaging data to high-fidelity geometric representations that offer high-resolution surface visualization [18]. The solution also necessitates the integration of advanced surface fitting methods that can facilitate the automatic generation of patient-specific models, which can be directly used for computational biomechanical analysis [19].
In this paper, we present a novel surface fitting framework that automatically generates patient-specific spline-based surface geometry of the tricuspid valve. The proposed method performs smooth and regular surface deformation of a template surface to fit the 3D point cloud obtained from echocardiographic image segmentation. We validate our framework using tricuspid valve surfaces obtained from simulation data, showing robustness toward different point cloud sampling and noise intensities. We then demonstrate the fitting results using point cloud data from real patients and the application of isogeometric analysis on the fitted surfaces  [20]. The rest of the article is organized as follows: In Section 2, we provide an overview of patient-specific computational modeling frameworks for heart valves. In Section 3, we explain the acquisition of echocardiographic imaging data obtained from patients and the generation of patient-specific point clouds from echocardiographic data through image segmentation. The proposed ValveFit framework is introduced in Section 4, and the surface fitting results along with a discussion on the performance are provided in Section 5. We also demonstrate that the surfaces generated by ValveFit are analysis-suitable by performing a simplified valve closure simulation. Finally, we provide concluding remarks and outline future directions in Section 6.
@@ 2 Related Work on Patient-Specific Modeling of Heart Valves
Heart valve surface fitting methods play a crucial role in the functional and structural assessment of the valves by providing accurate geometric representations from medical imaging data [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Several frameworks have been presented in the realm of patient-specific modeling based on surface fitting of point clouds obtained from image segmentation [37, 38]. For example, least squares optimization has been used to fit surfaces to point cloud segmentation data for the tricuspid valves [39]. These methods rely on good-quality image segmentation data to generate analysis-suitable meshes. If the point cloud data from images is sparse, noisy, and incomplete, these existing methods will likely overfit the data, resulting in inaccurate patient-specific valve models that would require considerable manual intervention to fix the models in order to achieve good mesh quality for analysis [40, 41]. Biomechanical models have been used to accurately fit the physiology of the heart valves [8, 42, 43, 44]. However, these methods do not impose stricter geometric regularization constraints to ensure good-quality surfaces during fast valve motion. On the other hand, machine learning-based approaches have been used to model the heart valve geometry directly from medical images, circumventing the image segmentation procedure [45]. However, the accuracy of these methods is dependent on large data from diverse populations, which is typically not available.
@@
  • A B-spline surface fitting framework using automatic differentiation, enabling faster B-spline surface evaluations and gradient computation.
@@
3 Image Acquisition and Segmentation
@@ 4.1 The Template Geometry
4.2 The Loss Function
The proposed ValveFit framework performs iterative deformation through the minimization of a loss function using gradient descent optimization. To ensure that the result of our surface fitting framework is an analysis-suitable surface approximating the patient-specific TV geometry [81, 82], we construct a loss function which quantifies the shape dissimilarity between the template surface and the target point cloud along with surface mesh regularity. The loss function ($\mathcal{L}$), which consists of a shape fidelity term and a surface regularization term is defined as
@@ 4.2 The Loss Function
where $\mathcal{E}_{\text{fid}}(\mathbf{S},\mathbf{Q})$ is the shape fidelity loss and $\mathcal{E}_{\text{reg}}(\mathbf{S})$ is the surface regularization energy. $\mathbf{S}$ denotes the template B-spline surface, and $\mathbf{Q}$ denotes the target point cloud.
4.2.1 Shape Fidelity Loss
The shape fidelity loss $\mathcal{E}_{\text{fid}}(\mathbf{S},\mathbf{Q})$ ensures that the fitted surface closely approximates the target point cloud. It is defined as
@@ 4.2.1 Shape Fidelity Loss
Here, $d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})$ is the one-sided Chamfer distance, $d_{\mathrm{HD}}(\mathbf{S},\mathbf{Q})$ is the Hausdorff distance, and $d_{\mathrm{a}}(\mathbf{A}_{\text{S}},\mathbf{A}_{\text{Q}})$ enforces the annulus constraint. The weights $w_{\text{CD}}$, $w_{\text{HD}}$, and $w_{\text{a}}$ control the influence of each term.
Chamfer Distance ($d_{\text{CD}}$). It measures the average discrepancy between two point clouds by calculating the mean nearest-neighbor distance in both directions. Given a template point set $\mathbf{S}$ and a target point set $\mathbf{Q}$, the symmetric Chamfer distance, $d^{sym}_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})$, is defined as
$d^{sym}_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})=\frac{1}{M}\sum_{k=1}^{M}\min_{l}\|\mathbf{s}_{k}-\mathbf{q}_{l}\|^{2}+\frac{1}{N}\sum_{l=1}^{N}\min_{k}\|\mathbf{q}_{l}-\mathbf{s}_{k}\|^{2}.$ (4)
Here $\mathbf{s}_{k}=\mathbf{S}(u_{k},v_{k})$, for k=1,,Mk=1,...,M, denote points sampled from the template surface $\mathbf{S}$, and $\mathbf{q}_{l}\in\mathbf{Q}$, for l=1,,Nl=1,...,N, denote points in the target point cloud $\mathbf{Q}$. The first term measures the mean nearest neighbor distance from the template surface points to the target point cloud ($\mathbf{S}\to\mathbf{Q}$), while the second term measures the mean nearest neighbor distance from the target point cloud to the template surface points ($\mathbf{Q}\to\mathbf{S}$). While the symmetric Chamfer distance is a comprehensive measure of bidirectional similarity, it can introduce artifacts when the target point cloud $\mathbf{Q}$ is sparse. The first term ($\mathbf{S}\to\mathbf{Q}$) forces every point on the dense template surface to align closely with one of the few sparse target points, which may cause the template surface to collapse or self-intersect in order to minimize this distance. To avoid these artifacts and ensure a more robust fit, we instead employ a one-sided Chamfer distance that considers only the second term, measuring the nearest neighbor distance from the target point cloud $\mathbf{Q}$ to the template surface $\mathbf{S}$. This is defined as
$d_{\mathrm{CD}}(\mathbf{S},\mathbf{Q})=\frac{1}{N}\sum_{l=1}^{N}\min_{k}\|\mathbf{q}_{l}-\mathbf{s}_{k}\|^{2}.$ (5)
Fig. 5 demonstrates the effect of the Chamfer distance weight $w_{\text{CD}}$. A weight that is too low leads to underfitting, where the surface fails to accurately fit the target point cloud, while a weight value that is too high causes overfitting and introduces artifacts such as surface folding.
Figure 5: Effect of the Chamfer distance weight, $w_{\text{CD}}$, on surface fitting accuracy. (a) With no contribution from the Chamfer distance term ($w_{\text{CD}}=0$), the surface fails to accurately fit to the target point cloud. (b) A moderate weight value ($w_{\text{CD}}=20$) yields a well-fitted surface that accurately approximates the target point cloud. (c) An excessive large weight value ($w_{\text{CD}}=500$) leads to overfitting, resulting in unnatural surface folding.
Hausdorff Distance ($d_{\text{HD}}$). The Hausdorff distance term ensures that the maximum deviation between the template surface and the target point cloud is minimized, thus ensuring that even the most distant points, often located near the boundaries, are closely approximated. It is defined as
$d_{\mathrm{HD}}(\mathbf{S},\mathbf{Q})=\max\left(\max_{k}\,\min_{l}\|\mathbf{s}_{k}-\mathbf{q}_{l}\|,\max_{l}\,\min_{k}\|\mathbf{q}_{l}-\mathbf{s}_{k}\|\right).$ (6)
The use of a one-sided Chamfer distance allows the template surface boundary to expand beyond the extent of the target point cloud without incurring a penalty, as it only enforces the proximity of the target points to the template surface. In contrast, the Hausdorff distance term explicitly penalizes the maximum deviation between the surface and the point cloud, thus constraining surface growth beyond the local neighborhood of the target point cloud. This effect is illustrated in Fig. 6 by varying the weight ($w_{\text{HD}}$) associated with the Hausdorff distance in Eq. (6), where an enlarged template surface is fitted to the target point cloud.
Figure 6: Effect of Hausdorff distance weight, $w_{\text{HD}}$, on constraining surface boundaries. (a) With $w_{\text{HD}}=0$, the right-bottom boundary of the surface extends significantly beyond the target point cloud. (b) A moderate weight value ($w_{\text{HD}}=1$) pulls the surface boundary closer, improving alignment with the target point cloud. (c) A higher weight value of $w_{\text{HD}}=10$ enforces tight adherence of the surface boundary to the boundary target points.
Annulus Constraint ($d_{\text{a}}$). The term $d_{\mathrm{a}}(\mathbf{A}_{\text{S}},\mathbf{A}_{\text{Q}})$ calculates the symmetric Chamfer distance (Eq. (4)) between the parametric boundary on the template surface closest to the annulus $\mathbf{A}_{\text{S}}:=\mathbf{S}(u=u_{0})$ and the segmented points corresponding to the annulus in the target point cloud $\mathbf{A}_{\text{Q}}$. The effect of this constraint is demonstrated in Fig. 7 by varying its corresponding weight, $w_{a}$. When the weight is zero (Fig. 7(a)), the annulus constraint is disabled. This results in a loose conformation that does not explicitly capture the target annulus points. As $w_{a}$ increases (Fig. 7(b)-(c)), this term enforces a better alignment, pulling the template surface boundary to achieve a tighter fit with the target annulus points.
Figure 7: Effect of the weight, $w_{a}$, for fitting to the target annulus points (red). (a) Weight of zero; (b) a moderate weight aligns the surface boundary with the target annulus points; and (c) a higher weight enforces a tighter conformation.
4.2.2 Surface Regularization Energy
@@
The surface regularization energy $\mathcal{E}_{\text{reg}}(\mathbf{S})$ imposes smoothness and regularity constraints on the deforming surface to obtain a smooth parameterization in the physical domain and prevent element distortions or self-intersections. It comprises three terms:
$\mathcal{E}_{\text{reg}}(\mathbf{S})=w_{\text{orth}}\,R_{\text{orth}}(\mathbf{S})+w_{\text{TPE}}\,R_{\text{TPE}}(\mathbf{S})+w_{\text{norm}}\,R_{\text{norm}}(\mathbf{S}).$ (7)
@@
The weights $w_{\text{orth}}$, $w_{\text{norm}}$, and $w_{\text{TPE}}$ allow us to control the influence of each of the above three terms in Eq. (7), allowing for adjustments based on specific mesh quality requirements.
Figure 8: The effect of varying the orthogonality weight, $w_{\text{orth}}$. (a) Distribution of the angle between tangent vectors, showing improved orthogonality with higher weight values. A visualization of the (b) surface elements and (c) fitted surface at $w_{\text{orth}}=0$ and $w_{\text{orth}}=20$.
Tangent Orthogonality Energy ($R_{\text{orth}}$). To enforce bijectivity during surface deformation, we need to avoid self-intersections and preserve surface mesh element quality throughout the transformation [83]. In particular, we ensure good surface mesh quality by penalizing extreme mesh distortion in each element. This is achieved by introducing an orthogonality regularization term, $R_{\text{orth}}(\mathbf{S})$, which ensures that the two parametric lines in the physical domain are locally orthogonal, thus maintaining surface regularity during deformation. This term is defined as
[…]
where $\mathbf{t}_{k}^{u}=\mathbf{t}^{u}(u_{k},v_{k})$ and $\mathbf{t}_{k}^{v}=\mathbf{t}^{v}(u_{k},v_{k})$ denote the tangent vectors evaluated at each sampled parametric point $(u_{k},v_{k})$, and $M$ is the number of sampled points in the parametric domain. Minimizing this energy results in an orthogonal parameterization that is free of local shearing. This angle preserving quality is critical for ensuring an analysis-suitable surface. By calculating the cosine of the angle between the local tangent vectors at each parametric point, $R_{\text{orth}}$ term penalizes deviations from right angles, thus avoiding excessive skewing or twisting of surface elements during deformation. The influence of this term is controlled by a weight parameter, $w_{\text{orth}}$, which introduces a trade-off between geometric regularity and data-fitting accuracy. As illustrated in Fig. 8, setting $w_{\text{orth}}$ to zero allows the surface to deform freely but can result in skewed mesh elements. Conversely, a higher weight value produces a smoother, more regular surface with a higher percentage of surface points with near locally orthogonal tangent vectors. $R_{\text{orth}}$ also ensures that the area of each mesh element remains positive during deformation, thus, also ensuring that the surface has no local degeneracies (it does not collapse to a point or line).
Figure 9: Effect of varying the Tangent Point Energy (TPE) parameters, $\alpha$ and $w_{\text{TPE}}$. (a) Distribution of TPE values across all surface point pairs for different $\alpha$ values. (b) TPE versus inter-point distance for different $\alpha$ values. (c) Fitted surface with $w_{\text{TPE}}=0$, resulting in self-intersection, and with $w_{\text{TPE}}=5.0$, where TPE effectively prevents collisions.
Tangent Point Energy ($R_{\text{TPE}}$). While the tangent orthogonality energy preserves local element shape quality, it does not prevent global self-intersections, such as collisions between leaflets during coaptation. To address this, we incorporate a tangent-point energy (TPE) term, $R_{\text{TPE}}(\mathbf{S})$, which introduces a repulsive potential that imposes an infinite energy barrier between any two surface points that approach each other [84, 85]. This term penalizes the most closest interaction between all points on the surface, a strategy shown empirically to be more effective at preventing self-collisions than averaging over all interactions. This term is defined as
$R_{\text{TPE}}(\mathbf{S})=\max_{k\neq l}\left(\frac{\left|\mathbf{n}_{k}\cdot(\mathbf{s}_{k}-\mathbf{s}_{l})\right|^{\alpha}}{\left\|\mathbf{s}_{k}-\mathbf{s}_{l}\right\|^{2\alpha}}\right),$ (9)
where $\mathbf{s}_{k}=\mathbf{S}(u_{k},v_{k})$ and $\mathbf{s}_{l}=\mathbf{S}(u_{l},v_{l})$ are any two distinct points on the surface, $\mathbf{n}_{k}$ is the unit normal at $\mathbf{s}_{k}$, and $\alpha$ controls the strength and decay rate of the repulsive interaction force.
Fig. 9 illustrates the influence of $\alpha$ and the weight $w_{\text{TPE}}$ on the TPE. $\alpha$ modulates the interaction range and sharpness of the repulsive potential. As shown in Fig. 9(b), higher values of $\alpha$ produce a steep, short-range repulsion, while lower values result in a gentler, long-range force. This behavior directly affects which point pairs are penalized; smaller $\alpha$ values distribute the repulsive energy across a broader set of pairs, whereas larger $\alpha$ values concentrate the penalty on a small subset of the closest set of pairs, as seen in Fig. 9(a). The accuracy of the TPE evaluation depends on the surface sampling strategy. Instead of carrying out adaptive quadrature [85], we adopt a uniform sampling approach for computational efficiency. This method has proven sufficient to prevent self-intersections for the valve geometries considered in this study. The parameter value $\alpha=4.0$ is selected as it provides an effective balance between the repulsive force strength and its interaction range. As seen in Fig. 9(a) $\alpha=4.0$ focuses the penalty on a moderate group of high-risk point pairs more effectively than $\alpha=2.0$, while avoiding the narrow and unstable energy landscape associated with higher $\alpha$ values. The resulting $\alpha=4.0$ curve in Fig. 9(b) exhibits a smoother and more progressive onset of the repulsive contact force, resulting in stable convergence during optimization. We thus use $\alpha=4.0$ for all the examples in Section 5. In Fig. 9(c), we showcase how by setting $w_{\text{TPE}}=5.0$, we can avoid leaflet collisions during coaptation by introducing the tangent point energy term.
Normal Deviation Energy ($R_{\text{norm}}$). The surface fitting process can be sensitive to noise in the target point cloud, often resulting in undesirable folding and self-intersections. To prevent this, we introduce a normal deviation energy term that reduces abrupt changes in the surface orientation. Specifically, by controlling the variation of the vertical ($z$) component of the unit normal vector $\mathbf{n}$, this term penalizes excessive folding, bending and self-intersections of the template surface. The normal deviation energy is defined as
$R_{\text{norm}}(\mathbf{S})=\max_{1\leq k\leq M}\left|\mathbf{n}_{k}^{z}-\bar{\mathbf{n}}^{z}\right|,$ (10)
where $\mathbf{n}_{k}^{z}$ denotes the $z$-component of the unit normal vector $\mathbf{n}$ evaluated at the parametric coordinate $(u_{k},v_{k})$, and $\bar{\mathbf{n}}^{z}$ is the mean value of the $z$-component of the normal vectors across all the surface points. By minimizing this term, abrupt changes in surface orientation are prevented, suppressing excessive folding of the surface that may arise from fitting noisy point clouds. As demonstrated in Fig. 10, increasing the weight $w_{\text{norm}}$ visibly regularizes the surface by enforcing a consistent alignment of the normal vectors and thus a more smoother geometry. Additionally, this term can be used as a soft constraint to control the shape of the surface, particularly in regions with sparse point cloud coverage, where data-driven guidance for fitting is difficult.
Figure 10: Effect of the normal deviation weight, $w_{\text{norm}}$, on surface smoothness and normal vector orientation. (a) Without regularization ($w_{\text{norm}}=0$), the normal vectors intersect each other, particularly in the bottom right corner, resulting in the folding of the surface. (b) A moderate weight ($w_{\text{norm}}=5.0$) reduces surface folding by enhancing consistent orientation of the normals. (c) A high weight ($w_{\text{norm}}=20.0$) produces a regular surface with a well-aligned field of normal vectors.
We can efficiently evaluate the tangent and normal vectors required for computing the terms in Eq. (7) by using the automatic differentiation capabilities of JAX [86]. We evaluate the partial derivatives of the surface coordinates ($x$, $y$, and $z$) with respect to the parametric coordinates $u$ and $v$ by evaluating the Jacobian matrix ($J$) of the B-spline surface parameterization as